the hypotenuse of a right triangle PQR is 16 feet long. the sine of angle P is 3/5, about how many feet long is QR ?
TRIANGLE
P
|.
|.....
|.......
|.........
|...........
|.............
|------------- R
Q

Answer by stanbon(75887)

the hypotenuse of a right triangle PQR is 16 feet long. the sine of angle P is 3/5, about how many feet long is QR ?
TRIANGLE
P
|.
|.....
|.......
|.........
|...........
|.............
|------------- R
Q
===============================
sine(P) = QR/PR = 3/5
But PR = 16
So QR = 16(3/5)
QR = 48/5 = 9 3/5 ft
===========================